OpenStudy (anonymous):

Can anyone do this??? The annual interest on a 19000 investment exceeds the interest earned on a 18000 investment by 204. the 19000 is invested at a .6% higher rater than the 18000. What is the interest rate of each investment?

OpenStudy (anonymous):

i messed this up somehow let me start again

OpenStudy (anonymous):

ok. i'm quite confused

OpenStudy (anonymous):

yeah you should be because i was being stupid. let me start over

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

put the rate of the 18000 investment at r the rate of the 19000 investment is r+.006 we have to work with numbers, not percents

OpenStudy (anonymous):

then you know that the return on 18000 is \[18000\times r\] and the return on 19000 is \[19000\times (r+.006)\] and you know that the second one is $204 more than the first, so \[18000\times r +204=19000(r+.006)\] now we can solve for r

OpenStudy (anonymous):

\[18000r+204=19000r + .006\times 19000\] \[18000r + 204=19000 r + 114\] \[204-114=19000r -18000r\] \[90=1000r\] \[r=\frac{90}{1000}=\frac{9}{100}=9\%\]

OpenStudy (anonymous):

love to know where they got 9% interest , i would go invest there

OpenStudy (anonymous):

i have to have the interest for both. do i add or subtract .6%

OpenStudy (anonymous):

?